# DPQuant: Efficient and Differentially-Private Model Training via Dynamic Quantization Scheduling

### v1

@@ A.3 Proof of Proposition 1
Let $q:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be an unbiased ($\mathbb{E}[q(\mathbf{x})]=\mathbf{x}$) and scale-invariant ($q(\lambda\mathbf{x})=\lambda\,q(\mathbf{x})$) quantizer whose outputs lie on a fixed finite grid. If $\mathbf{x}$ is drawn from an absolutely continuous distribution, then
$\operatorname{Var}\bigl{(}q(\mathbf{x})\bigr{)}\;=\;\Theta\bigl{(}\|\mathbf{x}\|_{\infty}^{2}\bigr{)}.$
@@
$\operatorname{Var}\bigl{(}q(\mathbf{x})\bigr{)}=M^{2}\operatorname{Var}\left(q(\mathbf{v})\right)\leq C\,M^{2}=C\,\|\mathbf{x}\|_{\infty}^{2}.$
@@
From this we conclude $\operatorname{Var}(q(\mathbf{x}))=\Theta(\|\mathbf{x}\|_{\infty}^{2})$, as desired. ∎
A.4 Justification of Privacy Guarantees (Theorem 2)
We issue a correction of the Proposition 2, listed below (with the difference highlighted in red).
@@ A.4 Justification of Privacy Guarantees (Theorem 2)
In step (3) of Algorithm 1, we clip the vector $\mathbf{R}$ to norm $\mathcal{C}$, to which independent Gaussian noise proportional to $\sigma^{2}\mathcal{C}^{2}$ is added to obtain $\hat{\mathbf{R}}$. This is equivalent to adding noise proportional to $\sigma^{2}$ when the sensitivity of $\hat{\mathbf{R}}$ is $1$ through a scaling argument.
Algorithm 1 ceases to access private data in $B$ after the computation of $\hat{\mathbf{R}}$, which results in all the following steps (i.e. updating privacy accountant and EMA) post-processing Dwork and Roth [2014] which does not impact the privacy consumption.
Furthermore, the privacy accounting step makes use of Opacus’ Yousefpour et al. [2022] privacy accountant, which assumes [^2] This assumption is stated in https://github.com/pytorch/opacus/blob/main/opacus/accountants/analysis/rdp.py the the noise scale $\sigma$ is proportional to the clipping constant (i.e. equivalent to adding a noise proportional to $\sigma^{2}\mathcal{C}^{2}$. ∎
A.5 Low Precision Simulation Setup
As FP4 hardware support is forthcoming, we employ the following simulation setup to emulate the effect of traitning under FP4. Notably, we quantize both inputs to the conv2d forward, wgrad, and dgrad operators as well as its output.
@@ A.5 Low Precision Simulation Setup
A.6 Theoretical Speedup Calculation
Due to the unavailabilty of accelerators and reliable software support for FP4, we instead rely on a performance model to estimate the theoretical throughputs of FP4 computation.
@@ A.6 Theoretical Speedup Calculation
We performed profiling on the models (on their respective datasets) stated in the paper, and we plot the runtime decomposition in Figure 8. Using this data, we can compute the amount of “overhead” (i.e. the time spent on operators which will not benefit from lower precision) for each model/dataset. This is tabulated in Table  7.
@@
A.7 Opacus Privacy Accounting
@@
A.8 Sampling of Quantized Layers
We outline the algorithm DPQuant uses to select layers to compute under quantization.

### v2

@@ A.10 Proof of Proposition 1
Let $q:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be an unbiased ($\mathbb{E}[q(\mathbf{x})]=\mathbf{x}$) and scale‐invariant ($q(\lambda\mathbf{x})=\lambda\,q(\mathbf{x})$) quantizer whose outputs lie on a fixed finite grid. If $\mathbf{x}$ is drawn from an absolutely continuous distribution, then
[…]
@@
[…]
@@
[…]
A.11 Justification of Privacy Guarantees (Theorem 2)
@@
In step  3 of Algorithm 1, we clip the vector $\mathbf{R}$ to norm $\mathcal{C}$, to which independent Gaussian noise proportional to $\sigma^{2}\mathcal{C}^{2}$ is added to obtain $\hat{\mathbf{R}}$. This is equivalent to adding noise proportional to $\sigma^{2}$ when the sensitivity of $\hat{\mathbf{R}}$ is $1$ through a scaling argument.
Algorithm 1 ceases to access private data in $B$ after the computation of $\hat{\mathbf{R}}$, which results in step  4 (updating the policy EMA) being post-processing Dwork and Roth (2014) which does not impact the privacy consumption.
Furthermore, the privacy accounting step makes use of Opacus’ Yousefpour et al. (2022) privacy accountant, which assumes [^2] This assumption is stated in https://github.com/pytorch/opacus/blob/main/opacus/accountants/analysis/rdp.py the noise scale $\sigma$ is proportional to the clipping constant (i.e. equivalent to adding a noise proportional to $\sigma^{2}\mathcal{C}^{2}$. ∎
A.12 Low Precision Simulation Setup
As FP4 hardware support is forthcoming, we employ the following simulation setup to emulate the effect of training under FP4. Notably, we quantize both inputs to the conv2d forward, wgrad, and dgrad operators as well as its output.
@@ A.12 Low Precision Simulation Setup
A.13 Theoretical Speedup Calculation
Due to the unavailability of accelerators and reliable software support for FP4, we instead rely on a performance model to estimate the theoretical throughputs of FP4 computation.
@@ A.13 Theoretical Speedup Calculation
We performed profiling on the models (on their respective datasets) stated in the paper, and we plot the runtime decomposition in Figure 8. Using this data, we can compute the amount of “overhead” (i.e. the time spent on operators which will not benefit from lower precision) for each model/dataset. This is tabulated in Table 14.
@@
A.14 Opacus Privacy Accounting
@@
A.15 Sampling of Quantized Layers
We outline the algorithm DPQuant uses to select layers to compute under quantization in Algorithm 2.
@@ A.15 Sampling of Quantized Layers
A.16 Remark: Applicability to Non-DP Training
The mechanisms in Sections 5.1 and 5.2 are general and could in principle apply to non-DP settings. However, prior work such as LUQ-FP4 (Chmiel et al., 2024) shows that FP4 non-DP training suffers negligible degradation, so DPQuant’s improvements would be marginal there. Our main novelty and motivation are DP-specific: we identify and analyze a setting where non-DP FP4 training shows negligible degradation, but DP + FP4 suffers substantial accuracy loss (Figure 1(a)), attributing this to the interaction between $\ell_{2}$-calibrated DP noise and $\ell_{\infty}$-driven quantization. Furthermore, our loss-sensitivity measures and selection rules are designed to be compatible with DP accounting and composition, relying only on quantities that are already DP-released or incorporable into the privacy accountant, which is not a concern in standard non-DP quantization work.
A.17 Remark: Vulnerability to Floating Point Attacks
Differential privacy implementations must carefully consider the vulnerabilities highlighted by Mironov (2012). Mironov identified that the floating-point implementation of noise sampling for mechanisms such as Laplacian or Gaussian introduces a “porous” distribution that lacks translation invariance. This issue is prevalent in both fp64 and fp32 arithmetic.
To ensure robustness against this vulnerability, our method has been meticulously designed. The critical step of noise addition in our framework occurs under standard conditions, prior to the application of our novel quantization technique. The process is as follows:
  1. 1.
    Gradients are maintained in full fp32 precision.
  2. 2.
    Noise is sampled and added to these fp32 gradients, also in fp32 precision. Only after the noisy gradient is computed is it quantized for use in the forward/backward pass of select layers.
  3. 3.
    Thus, the noise injection process maintains a vulnerability profile identical to that of standard DP-SGD implemented in fp32. The use of lower-precision representations for computation does not alter or exacerbate the known properties of the initial noise addition.
Additionally, our method is fully compatible with established defenses against this vulnerability. The ’snapping mechanism’ proposed by Mironov, a post-processing step applied directly to the noisy output, would be applied to the full-precision fp32 gradients immediately after noise addition and before quantization in our pipeline.